# Ctrl-World: A Controllable Generative World Model for Robot Manipulation

### v2

@@ 2 Related Works
Action-Conditioned World Models. Although pretrained video models are powerful, they are often only conditioned on high-level language instructions. Nonetheless, some prior works have explored using action-conditioned predictive models, both in low-dimensional state spaces (Nagabandi et al., 2020) and with image observations Hafner et al. (2019; 2020); Hansen et al. (2022); Wu et al. (2023); Oh et al. (2015). Many of these approaches learn task-specific models (Hafner et al., 2019), while we focus on training generalist, multi-task world models. Building on early works (Finn & Levine, 2017; Ebert et al., 2018; Xie et al., 2019; Dasari et al., 2019; Yang et al., 2023; Wu et al., 2024) as well as more recent approaches that leverage diffusion (Quevedo et al., 2025; Chen et al., 2024; Ball et al., 2025; Gao et al., 2025) and frame-level action conditioning (Zhu et al., 2024), we propose a model that incorporates multi-view prediction, long-horizon temporal coherence, and fine-grained controllability. Our experiments show that these capabilities enable effective evaluation and improvement of state-of-the-art generalist VLA policies.
@@ 4.1 Learning World Model Ctrl-World
Multi-View Joint Predictions. State-of-the-art VLA models often rely on multiple third-person cameras for global context and wrist-mounted cameras for precise interactions (Intelligence et al., 2025; Liu et al., 2024; 2025). To match this, the world model must generate spatially consistent predictions across all views at each step. Prior work has shown that feed-forward transformers can effectively capture spatial relationships between multi-view cameras in a scalable manner (Wang et al., 2025). Following prior work, we concatenate the $N$ input images—each containing $H\times W$ tokens—along the token dimension and jointly predict all views $o_{t:t+H}$. In experiments, we find multi-view joint prediction also improves consistency and substantially reduces hallucinations.
@@
Frame-level Action Conditioning. The pretrained video model conditions only on text and image, which limits its control precision. To enable full controllability, we additionally condition the model on the action sequence $a_{t+1:t+H}]$ output by the policy. We also transform each action sequence into Cartesian-space robot arm poses $a^{\prime}_{t+1:t+H}]$ and concatenate with past poses $q_{t-km},...,q_{t-m},q_{t}]$. Frame-wise cross-attention is then applied within the spatial transformer, allowing the visual tokens of each frame to attend to its associated pose embedding. For history frames, this pose corresponds to $q_{t-km},...,q_{t-m},q_{t}]$, while for future frames, it corresponds to $a^{\prime}_{t+1:t+H}]$.
@@ 5.1 Experiment Setups
Third-view
Camera
WPE 20.33 0.772 0.131 25.50 156.4
IRASim 21.36 0.774 0.117 26.46 138.1
Ctrl-World-Third-View 21.27 0.793 0.110 23.47 127.5
@@
Evaluated
Camera
Method Computation-based Model-based
@@
Third-view
Camera
Ctrl-World 23.56 0.828 0.091 25.00 97.4
@@
Wrist-view
Camera
Ctrl-World 19.18 0.665 0.252 25.78 127.1
@@ 5.2 World Model Quality Analysis
Baselines and Evaluation Matrices. We quantitatively compare our model, Ctrl-World, against two prior action-conditioned world models: World-model-based Policy Evaluation (WPE) (Quevedo et al., 2025) and IRASim (Zhu et al., 2024). Since these models only predict from a single third-person camera view, we train a single-view version, Ctrl-World-third-view, which only inputs and predicts on a single third-person camera for a fair comparison. For evaluation, we hold out 2% of the trajectories as a validation set and randomly sample 256 video clips, each 10 s in length. During rollouts, the world model receives 15-step action chunks (corresponding to 1 s) and autoregressively predicts the next frames for 10 steps, producing 10 s-long trajectories. We then compare the predicted videos against ground truth using both computational (PSNR (Hore & Ziou, 2010) and SSIM (Wang et al., 2004)) and model-based metrics (LPIPS (Zhang et al., 2018), FID (Heusel et al., 2017), and FVD (Unterthiner et al., 2018)).
@@ 5.3 World Model for Policy Evaluation
In this section, we evaluate whether Ctrl-World can be used to evaluate the instruction-following ability of generalist robot policies and accurately reflect their performance rankings in the real world. We set up our own DROID platform and randomly place two third-person cameras around the workspace. Similar to how prior works have seen DROID policies generalize to new setups (Pertsch et al., 2025), we find that Ctrl-World, pretrained solely on the open-sourced DROID dataset, can make accurate future predictions zero-shot in our newly configured scene with novel camera placements.
@@ 6 Conclusion
7 ACKNOWLEDGMENTS
We would like to thank Ajay Sridhar and Ke Wang for their help when setting up the hardware. We also thank Moo Jin Kim, James Qian, and Yuejiang Liu for insightful discussions on the ideas. This work was supported by the Toyota Research Institute (TRI), ONR grant N00014-22-1-2621, and the Robotics and AI Institute.

### v3

@@ 2 Related Works
Action-Conditioned World Models. Although pretrained video models are powerful, they are often only conditioned on high-level language instructions. Nonetheless, some prior works have explored using action-conditioned predictive models, both in low-dimensional state spaces (Nagabandi et al., 2020) and with image observations Hafner et al. (2019; 2020); Hansen et al. (2022); Wu et al. (2023); Oh et al. (2015); Huang et al. (2025). Many of these approaches learn task-specific models (Hafner et al., 2019), while we focus on training generalist, multi-task world models. Building on early works (Finn and Levine, 2017; Ebert et al., 2018; Xie et al., 2019; Dasari et al., 2019; Yang et al., 2023; Wu et al., 2024) as well as more recent approaches that leverage diffusion (Quevedo et al., 2025; Chen et al., 2024; Ball et al., 2025; Gao et al., 2025; Ren et al., 2025; Hafner et al., 2025) and frame-level action conditioning (Zhu et al., 2024), we propose a model that incorporates multi-view prediction, long-horizon temporal coherence, and fine-grained controllability. Our experiments show that these capabilities enable effective evaluation and improvement of state-of-the-art generalist VLA policies.
@@ 4.1 Learning World Model Ctrl-World
Multi-View Joint Predictions. State-of-the-art VLA models often rely on multiple third-person cameras for global context and wrist-mounted cameras for precise interactions (Intelligence et al., 2025; Liu et al., 2024; 2025). To match this, the world model must generate spatially consistent predictions across all views at each step (Jiang et al., 2025). Prior work has shown that feed-forward transformers can effectively capture spatial relationships between multi-view cameras in a scalable manner (Wang et al., 2025). Following prior work, we concatenate the $N$ input images—each containing $H\times W$ tokens—along the token dimension and jointly predict all views $o_{t:t+H}$. In experiments, we find multi-view joint prediction also improves consistency and substantially reduces hallucinations.
@@
Frame-level Action Conditioning. The pretrained video model conditions only on text and image, which limits its control precision. To enable full controllability, we additionally condition the model on the action sequence $a_{t+1:t+H}]$ output by the policy. We also transform each action sequence into Cartesian-space robot arm poses $a^{\prime}_{t+1:t+H}]$ and concatenate with past poses $q_{t-km},...,q_{t-m},q_{t}]$. Frame-wise cross-attention (Zhu et al., 2024; He et al., 2025) is then applied within the spatial transformer, allowing the visual tokens of each frame to attend to its associated pose embedding. For history frames, this pose corresponds to $q_{t-km},...,q_{t-m},q_{t}]$, while for future frames, it corresponds to $a^{\prime}_{t+1:t+H}]$.
@@ 5.1 Experiment Setups
Third-view Camera WPE-Single-View 20.33 0.772 0.131 25.50 156.4
WPE-Multiview 21.17 - - - 147.1
IRASim-Single-View 21.36 0.774 0.117 26.46 138.1
IRASim-Multiview 20.21 - - - 165.4
Ctrl-World-Single-View 21.27 0.793 0.110 23.47 127.5
@@
Evaluated Camera Method Computation-based Model-based
@@
Third-view Camera Ctrl-World 23.56 0.828 0.091 25.00 97.4
@@
Wrist-view Camera Ctrl-World 19.18 0.665 0.252 25.78 127.1
@@ 5.2 World Model Quality Analysis
Baselines and Evaluation Matrices. We quantitatively compare our model, Ctrl-World, against two prior action-conditioned world models: World-model-based Policy Evaluation (WPE) (Quevedo et al., 2025) and IRASim (Zhu et al., 2024). Since these models only predict from a single third-person camera view, we train a single-view version, Ctrl-World-third-view, which only inputs and predicts on a single third-person camera for a fair comparison. For evaluation, we hold out 2% of the trajectories as a validation set and randomly sample 256 video clips, each 10 s in length. During rollouts, the world model receives 15-step action chunks (corresponding to 1 s) and autoregressively predicts the next frames for 10 steps, producing 10 s-long trajectories. We then compare the predicted videos against ground truth using both computational (PSNR (Hore and Ziou, 2010) and SSIM (Wang et al., 2004)) and model-based metrics (LPIPS (Zhang et al., 2018), FID (Heusel et al., 2017), and FVD (Unterthiner et al., 2018)).
@@ 5.3 World Model for Policy Evaluation
In this section, we evaluate whether Ctrl-World can be used to evaluate the instruction-following ability of generalist robot policies and accurately reflect their performance rankings in the real world (Li et al., 2024). We set up our own DROID platform and randomly place two third-person cameras around the workspace. Similar to how prior works have seen DROID policies generalize to new setups (Pertsch et al., 2025), we find that Ctrl-World, pretrained solely on the open-sourced DROID dataset, can make accurate future predictions zero-shot in our newly configured scene with novel camera placements.